# Quantum Speedups in Regret Analysis of Infinite Horizon Average-Reward Markov Decision Processes

### v2

@@ Abstract
I Introduction
Quantum Machine Learning (QML) has garnered remarkable interest in contemporary research, predominantly attributed to the pronounced speedups achievable with quantum computers as opposed to their classical analogs [1, 2]. The fundamental edge of quantum computing arises from the unique nature of its fundamental computing element, termed a qubit, which can exist simultaneously in states of 0 and 1, unlike classical bits that are restricted to either 0 or 1. This inherent distinction underpins the exponential advancements that quantum computers bring to specific computational tasks, surpassing the capabilities of classical computers.
In the realm of Reinforcement Learning (RL), an agent embarks on the task of finding an efficient policy for Markov Decision Process (MDP) environment through repetitive interactions [3]. RL has found notable success in diverse applications, including but not limited to autonomous driving, ridesharing platforms, online ad recommendation systems, and proficient gameplay agents [4, 5, 6, 7]. Most of these setups require decision making over infinite horizon, with an objective of average rewards. This setup has been studied in classical reinforcement learning [8, 9, 10], where $\tilde{\mathcal{O}}(\sqrt{T})$ regret across $T$ rounds has been found. This paper embarks on an inquiry into the potential utilization of quantum statistical estimation techniques to augment the theoretical convergence speeds of tabular RL algorithms within the context of infinite horizon learning settings.
A diverse array of quantum statistical estimation primitives has emerged, showcasing significant enhancements in convergence speeds and gaining traction within Quantum Machine Learning (QML) frameworks [11, 12, 13]. Notably, [14] introduces a quantum mean estimation algorithm that yields a quadratic acceleration in sample complexity when contrasted with classical mean estimation techniques. In this work, we emphasize the adoption of this particular mean estimation method as a pivotal component of our methodology. It strategically refines the signals garnered by the RL agent during its interaction with the enigmatic quantum-classical hybrid MDP environment.
@@ I Introduction
    Our paper introduces a pioneering contribution in the form of Q-UCRL, an innovative model-based, infinite horizon, optimism-driven quantum Reinforcement Learning (QRL) algorithm. Drawing inspiration from classical predecessors like the UCRL and UC-CURL algorithms [8, 9, 15], Q-UCRL is designed to seamlessly weave together an agent’s optimistic policy acquisition and an adept quantum mean estimator.
@@
    Employing meticulous theoretical analysis, we show that the Q-UCRL algorithm achieves an exponential improvement in the regret analysis. Specifically, it attains a regret bound of $\tilde{\mathcal{O}}(1)$, a significant improvement from the classical benchmark of $\tilde{\mathcal{O}}(\sqrt{T})$ across $T$ online rounds. The analysis is based on the novel quantum Bellman error based analysis (introduced in classical RL in [9] for optimism based algorithm), where the difference between the performance of a policy on two different MDPs is bounded by long-term averaged Bellman error and the quantum mean estimation is used for improved Bellman error.
@@ II Related Work
Infinite Horizon Reinforcement Learning: Regret analysis of infinite horizon RL in the classical setting has been widely studied with average reward objective. A prominent principle that underpins algorithms tailored for this scenario is the concept of “optimism in the face of uncertainty” (OFU). In this approach, the RL agent nurtures optimistic estimations of value functions and, during online iterations, selects policies aligned with the highest value estimates [10, 8, 9]. Additionally, it’s noteworthy to acknowledge that several methodologies are rooted in the realm of posterior sampling, where the RL agent samples an MDP from a Bayesian Distribution and subsequently enacts the optimal policy [16, 17, 18, 15]. In our study, we embrace the OFU-based algorithmic framework introduced in [9], and we extend its scope to an augmented landscape where the RL agent gains access to supplementary quantum information. Furthermore, we render a mathematical characterization of regret, revealing a $\tilde{\mathcal{O}}(1)$ bound, which in turn underscores the merits of astutely processing the quantum signals within our framework.
@@
Quantum Mean Estimation: The realm of mean estimation revolves around the identification of the average value of samples stemming from an unspecified distribution. Of paramount importance is the revelation that quantum mean estimators yield a quadratic enhancement when juxtaposed against their classical counterparts [19, 14]. The key reason for this improvement is based on the quantum amplitude amplification, which allows for suppressing certain quantum states w.r.t. the states that are desired to be extracted [11]. In Appendix B, we present a discussion around Quantum Amplitude Estimation and its applications in QML.
@@
In the following, we introduce the definition of key elements and results pertaining to quantum mean estimation that are critical to our setup and analysis. First, we present the definition of a classical random variable and its extension to a quantum random variable (q.r.v.).
Definition 1 (Random Variable).
@@
A finite random variable can be represented as $X:\Omega\to E$ for some probability space $(\Omega,P)$, where $\Omega$ is a finite sample set, $P:\Omega\to[0,1]$ is a probability mass function and $E\subset\mathbb{R}$ is the support of $X$. $(\Omega,P)$ is frequently omitted when referring to the random variable $X$.
Definition 2 (Quantum Random Variable).
A q-variable is a tuple $(\mathcal{H},U,M)$ where $\mathcal{H}$ is a finite-dimensional Hilbert space, $U$ is a unitary transformation on $\mathcal{H}$, and $M=\{M_{x}\}_{x\in E}$ is a projective measurement on $\mathcal{H}$ indexed by a finite set $E\subset\mathbb{R}$. Given a random variable $X$ on a probability space $(\Omega,P)$, we say that a q-variable $(\mathcal{H},U,M)$ generates $X$ when,
  • (1)
    $\mathcal{H}$ is a finite-dimensional Hilbert space with some basis $\{|\omega\rangle\}_{\omega\in\Omega}$ indexed by $\Omega$.
  • (2)
    $U$ is a unitary transformation on $\mathcal{H}$ such that $U|\mathbf{0}\rangle=\sum_{\omega\in\Omega}\sqrt{P(\omega)}|\omega\rangle$.
  • (3)
    $M=\{M_{x}\}_{x}$ is the projective measurement on $\mathcal{H}$ defined by $M_{x}=\sum_{\omega:X(\omega)=x}|\omega\rangle\langle\omega|$.
With the assumption that a quantum evaluation oracle can be accessed and comparison and rotation oracles can be efficiently implemented [14], a quantum experiment is defined as the process of applying any of the unitaries $U$, comparisons, rotations, their inverses or their controlled versions, or performing a measurement according to $M$.
Without loss of generality, we call random variable $X$ as a q.r.v. if it is generated by some q-variable $(\mathcal{H},U,M)$. Furthermore, [20] draws the connection between a random variable and q quantum random variable by introducing the quantum sampling oracle.
Definition 3 (Quantum sampling oracle).
For a discrete random variable $X$ with sample space $\mathcal{X}$, its quantum sampling oracle $O_{X}$ is defined as
$O_{X}|0\rangle\rightarrow\sum_{\bf{x}\in\mathcal{X}}\sqrt{P_{X}(\bf{x})}|\bf{x}\rangle\otimes|\rm{garbage}(\bf{x})\rangle$ (1)
where $P_{X}(\cdot)$ represents the probability mass function of $X$, and the last quantum register in the output represents possible emerging garbage quantum states during the implementations.
With the access to the quantum sampling oracle, a finite space discrete random variable $X$ could be treated as a quantum random variable with the quantum sampling oracle being the unitary transformation, and each query of the quantum sampling oracle is equivalent to one quantum experiment.
@@
Next, we present a key quantum mean estimation result that will be carefully employed in our algorithmic framework to utilize the q.r.v.s collected by the RL agent. One of the crucial aspects of Lemma 1 is that quantum mean estimation converges at the rate $\mathcal{O}(\frac{1}{n})$ as opposed to classical benchmark convergence rate $\mathcal{O}(\frac{1}{\sqrt{n}})$ for $n$ number of samples, therefore estimation efficiency quadratically.
Lemma 1 (Sub-Gaussian estimator [14]).
Let $X$ be a q.r.v. with mean $\mu$ and variance $\sigma^{2}$. Given $n$ i.i.d. samples of q.r.v. $X$ and a real $\delta\in(0,1)$ such that $n>\log(1/\delta)$, a quantum algorithm SubGaussEst$(X,n,\delta)$ (please refer to algorithm 2 in [14]) outputs a mean estimate $\hat{\mu}$ such that,
@@
$\mathbb{P}\left[|\hat{\mu}-\mu|\leq\frac{\sigma\log(1/\delta)}{n}\right]\geq 1-\delta.$ (2)
The algorithm performs $O(n\log^{3/2}(n)\log\log(n))$ quantum experiments.
@@
Quantum Reinforcement Learning: Within the realm of QRL, a prominent strand of previous research showcases exponential enhancements in regret through amplitude amplification-based methodologies applied to the Quantum Multi-Armed Bandits (Q-MAB) problem [21, 22]. Nevertheless, the theoretical underpinnings of the aforementioned Q-MAB approaches do not seamlessly extend to the QRL context since there is no state evolution in bandits.
@@
A recent surge of research interest has been directed towards Quantum Reinforcement Learning (QRL) [23, 24, 25]. It is noteworthy to emphasize that the aforementioned studies do not provide an exhaustive mathematical characterization of regret. [26] delves into QRL within the infinite horizon Markov Decision Process (MDP) learning framework, yet their focus rests on a discounted reward setup, where advancements in sample complexity are demonstrated through the lens of generative model assumptions. In contrast, our work aligns itself with the agent-environment Quantum hybrid interaction framework established in [27], which solely explores QRL within the finite horizon MDP setting. This is the first paper to study infinite horizon MDP with average reward objective.
III Problem Formulation
@@
We consider the problem of Reinforcement Learning (RL) in an infinite horizon Markov Decision Process characterized by $\mathcal{M}\triangleq(\mathcal{S},\mathcal{A},P,r,D)$, wherein $\mathcal{S}$ and $\mathcal{A}$ represent finite collections of states and actions respectively, pertaining to RL agent’s interaction with the unknown MDP environment. $P(s^{\prime}|s,a)\in[0,1]$ denotes the transition probability for next state $s^{\prime}\in\mathcal{S}$ for a given pair of previous state and RL agent’s action, i.e., $(s,a)\in\mathcal{S}\times\mathcal{A}$. Further, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ represents the reward collected by the RL agent for state-action pair $(s,a)$. In the following, we first present the additional quantum transition oracle that the agent could access during its interaction with the unknown MDP environment at every RL round.
@@
Quantum Computing Basics: In this section, we give a brief introduction to the concepts that are most related to our work based on [28] and [29]. Consider an $m$ dimensional Hilbert space $\mathbb{C}^{m}$, a quantum state $|x\rangle=(x_{1},...,x_{m})^{T}$ can be seen as a vector inside the Hilbert space with $\sum_{i}|x_{i}|^{2}=1$. Furthermore, for a finite set with $m$ elements $\xi=\{\xi_{1},...,\xi_{m}\}$, we can assign each $\xi_{i}\in\xi$ with a member of an orthonormal basis of the Hilbert space $\mathbb{C}^{m}$ by
@@
Quantum transition oracle: We utilize the concepts and notations of quantum computing in [21, 30, 27, 31] to construct the quantum sampling oracle of RL environments and capture agent’s interaction with the unknown MDP environment. We now formulate the equivalent quantum-accessible RL environments for our classical MDP $\mathcal{M}$. For an agent at step $t$ in state $s_{t}$ and with action $a_{t}$, we construct the quantum sampling oracles for indicator random variables $\mathbbm{1}[s_{t+1}=s_{i}|s_{t},a_{t}]$ for all $s_{i}\in\mathcal{S}$. Specifically, suppose there are two Hilbert spaces $\mathcal{\bar{S}}=\mathbb{C}^{|\mathcal{S}|}$ and $\mathcal{\bar{A}}=\mathbb{C}^{|\mathcal{A}|}$ containing the superpositions of the classical states and actions. We represent the computational bases for $\mathcal{\bar{S}}$ and $\mathcal{\bar{A}}$ as $\{|s\rangle\}_{s\in\mathcal{S}}$ and $\{|a\rangle\}_{a\in\mathcal{A}}$. We assume that we can implement the quantum sampling oracle $\mathcal{P}_{s_{i}}$ of the MDP’s transitions for all $s_{i}\in\mathcal{S}$ as follows:
  • The transition oracle $\mathcal{P}_{s_{i}}$ which returns the superposition over 00 and $1$ according to $\mathbbm{1}[s^{\prime}=s_{i}|s,a]$, the probability of the next state being $s_{i}$ given the current state $|s\rangle$ and action $|a\rangle$ is defined as:
    $\mathcal{P}_{s_{i}}:|s\rangle\otimes|a\rangle\otimes|0\rangle\rightarrow|\bm{\psi}_{s_{i}}\rangle_{(s,a)}$
    Where $|\bm{\psi}_{s_{i}}\rangle_{(s,a)}=|s\rangle\otimes|a\rangle\otimes[\sqrt{P(s^{\prime}=s_{i}|s,a)}|1\rangle+\sqrt{1-P(s^{\prime}=s_{i}|s,a)|}0\rangle]$.
@@
As described in Fig 1, at step $t$, the agent plays an action $a_{t}$ based on current state $s_{t}$ according to its policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$. Consequently, the unknown MDP Q-environment shares the next state information and reward i.e., $\{s_{t+1},r_{t}\}$. Additionally, the quantum transition oracle first encodes $s_{t}$ and $a_{t}$ into the corresponding basis of the quantum Hilbert spaces, producing basis $|s_{t}\rangle$ and $|a_{t}\rangle$, then it creates a collection of superpositioned quantum states $\bm{\psi}^{t}=\{|\bm{\psi}_{s_{i}}\rangle_{(s_{t},a_{t})}\}_{s_{i}\in\ \mathcal{S}}$. Each quantum state $\bm{\psi}^{t}_{i}=|\bm{\psi}_{s_{i}}\rangle_{(s_{t},a_{t})}$ obtained from the quantum transition oracle $\mathcal{P}_{s_{i}}$ is quantum query of the indicator $\mathbbm{1}[s^{\prime}=s_{i}|s_{t},a_{t}]$, and thus can be used for the Quantum Mean Estimator SubGaussEst outlined in Lemma 1 which in turn improves transition probability estimation and leads to exponentially better regret.
RL regret formulation: Consider the RL agent following some policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ on the unknown MDP Q-environment $\mathcal{M}$. Then, as a result of playing policy $\pi$, the agent’s long term average reward is given by:
@@
Then, according to [32], $\Gamma_{\pi}^{P}$ can be represented as the following:
@@
wherein, $V_{\pi}^{P}(\cdot;\cdot)$ is the discounted cumulative average reward by following policy $\pi$, $\rho_{\pi}^{P}$ is the steady-state occupancy measure (i.e., $\rho_{\pi}^{P}(s,a)\in[0,1],~{}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\rho_{\pi}^{P}(s,a)=1$). In the following, we introduce key assumptions crucial from the perspective of our algorithmic framework and its analysis. In this context, we introduce the notations $\{P_{\pi,s}^{t},T_{\pi,~{}s\rightarrow s^{\prime}}\}$ to denote the $t$ time step probability distribution obtained by playing policy $\pi$ on the unknown MDP $\mathcal{M}$ starting from some arbitrary state $s\in\mathcal{S}$; and, the actual time steps to reach $s^{\prime}$ from $s$ by playing policy $\pi$ respectively. With these notations, we introduce our first assumption pertaining to ergodicity of MDP $\mathcal{M}$.
@@
We emphasize that Assumption 2 is a commonly used assumption in RL algorithm development owing to the fact that in most setups rewards are constructed according to the underlying problem and is known beforehand [33, 34]. Next, we define the cumulative regret accumulated by the agent in expectation across $T$ time steps as follows:

### v3

@@ Abstract
1 Introduction
Quantum Machine Learning (QML) has garnered remarkable interest in contemporary research, predominantly attributed to the pronounced speedups achievable with quantum computers as opposed to their classical analogs (?, ?). The fundamental edge of quantum computing arises from the unique nature of its fundamental computing element, termed a qubit, which can exist simultaneously in states of 0 and 1, unlike classical bits that are restricted to either 0 or 1. This inherent distinction underpins the exponential advancements that quantum computers bring to specific computational tasks, surpassing the capabilities of classical computers.
In the realm of Reinforcement Learning (RL), an agent embarks on the task of finding an efficient policy for Markov Decision Process (MDP) environment through repetitive interactions (?). RL has found notable success in diverse applications, including but not limited to autonomous driving, ridesharing platforms, online ad recommendation systems, and proficient gameplay agents (?, ?, ?, ?). Most of these setups require decision making over infinite horizon, with an objective of average rewards. This setup has been studied in classical reinforcement learning (?, ?, ?), where $\tilde{\mathcal{O}}(\sqrt{T})$ regret across $T$ rounds has been found meeting the theoretical lower bound (?). This paper embarks on an inquiry into the potential utilization of quantum statistical estimation techniques to augment the theoretical convergence speeds of tabular RL algorithms within the context of infinite horizon learning settings.
A diverse array of quantum statistical estimation primitives has emerged, showcasing significant enhancements in convergence speeds and gaining traction within Quantum Machine Learning (QML) frameworks (?, ?, ?). Notably, (?) introduces a quantum mean estimation algorithm that yields a quadratic acceleration in sample complexity when contrasted with classical mean estimation techniques. In this work, we emphasize the adoption of this particular mean estimation method as a pivotal component of our methodology. It strategically refines the signals garnered by the RL agent during its interaction with the enigmatic quantum-classical hybrid MDP environment.
@@ 1 Introduction
    Our paper introduces a pioneering contribution in the form of Quantum-UCRL (Q-UCRL), an innovative model-based, infinite horizon, optimism-driven quantum Reinforcement Learning (QRL) algorithm. Drawing inspiration from classical predecessors like the UCRL and UC-CURL algorithms (?, ?, ?), Q-UCRL is designed to seamlessly weave together an agent’s optimistic policy acquisition and an adept quantum mean estimator.
@@
    Employing meticulous theoretical analysis, we show that the Q-UCRL algorithm achieves an exponential improvement in the regret analysis. Specifically, it attains a regret bound of $\tilde{\mathcal{O}}(1)$, breaking through the theoretical classical lower bound of $\Omega(\sqrt{T})$ across $T$ online rounds. The analysis is based on the novel quantum Bellman error based analysis (introduced in classical RL in (?) for optimism based algorithm), where the difference between the performance of a policy on two different MDPs is bounded by long-term averaged Bellman error and the quantum mean estimation is used for improved Bellman error. Further, the analysis avoids dependence on martingale concentration bounds and incorporates the phenomenon of state collapse following measurements, a crucial aspect in quantum computing analysis.
@@ 2 Related Work
Infinite Horizon Reinforcement Learning: Regret analysis of infinite horizon RL in the classical setting has been widely studied with average reward objective. A prominent principle that underpins algorithms tailored for this scenario is the concept of “optimism in the face of uncertainty” (OFU). In this approach, the RL agent nurtures optimistic estimations of value functions and, during online iterations, selects policies aligned with the highest value estimates (?, ?, ?). Additionally, it’s noteworthy to acknowledge that several methodologies are rooted in the realm of posterior sampling, where the RL agent samples an MDP from a Bayesian Distribution and subsequently enacts the optimal policy (?, ?, ?, ?). In our study, we embrace the OFU-based algorithmic framework introduced in (?), and we extend its scope to an augmented landscape where the RL agent gains access to supplementary quantum information. Furthermore, we render a mathematical characterization of regret, revealing a $\tilde{\mathcal{O}}(1)$ bound, which in turn underscores the merits of astutely processing the quantum signals within our framework.
@@
Quantum Mean Estimation: The realm of mean estimation revolves around the identification of the average value of samples stemming from an unspecified distribution. Of paramount importance is the revelation that quantum mean estimators yield a quadratic enhancement when juxtaposed against their classical counterparts (?, ?). The key reason for this improvement is based on the quantum amplitude amplification, which allows for suppressing certain quantum states w.r.t. the states that are desired to be extracted (?). In Appendix B, we present a discussion around Quantum Amplitude Estimation and its applications in QML.
@@
In the following, we introduce the definition of key elements and results pertaining to quantum mean estimation that are critical to our setup and analysis. First, we present the definition of a classical random variable and the quantum sampling oracle for performing quantum experiments.
Definition 1 (Random Variable, Definition 2.2 of (?)).
@@
A finite random variable can be represented as $X:\Omega\to E$ for some probability space $(\Omega,\mathbb{P})$, where $\Omega$ is a finite sample set, $\mathbb{P}:\Omega\to[0,1]$ is a probability mass function and $E\subset\mathbb{R}$ is the support of $X$. $(\Omega,\mathbb{P})$ is frequently omitted when referring to the random variable $X$.
To perform quantum mean estimation, we provide the definition of quantum experiment. This is analogous to the classical random experiments
Definition 2 (Quantum Experiment).
Consider a random variable $X$ on a probability space $(\Omega,2^{\Omega},\mathbb{P})$. Let $\mathcal{H}_{\Omega}$ be a Hilbert space with basis states $\{\lvert\omega\rangle\}_{\omega\in\Omega}$ and fix a unitary $\mathcal{U}_{\mathbb{P}}$ acting on $\mathcal{H}_{\Omega}$ such that
$\mathcal{U}_{\mathbb{P}}:\lvert 0\rangle\mapsto\sum_{\omega\in\Omega}\sqrt{\mathbb{P}(\omega)}\lvert\omega\rangle$
assuming $0\in\Omega$. We define a quantum experiment as the process of applying the unitary $\mathcal{U}_{\mathbb{P}}$ or its inverse $\mathcal{U}_{\mathbb{P}}^{-1}$ on any state in $\mathcal{H}_{\Omega}$.
The unitiary $\mathcal{U}_{\mathbb{P}}$ provides the ability to query samples of the random variable in superpositioned states, which is the essence for speedups in quantum algorithms. To perform quantum mean estimation for values of random variables (?, ?, ?), an additional quantum evaluation oracle would be needed.
Definition 3 (Quantum Evaluation Oracle).
Consider a finite random variable $X:\Omega\to E$ on a probability space $(\Omega,2^{\Omega},\mathbb{P})$. Let $\mathcal{H}_{\Omega}$ and $\mathcal{H}_{E}$ be two Hilbert spaces with basis states $\{|\omega\rangle\}_{\omega\in\Omega}$ and $\{|x\rangle\}_{x\in E}$ respectively. We say that a unitary $\mathcal{U}_{X}$ acting on $\mathcal{H}_{\Omega}\otimes\mathcal{H}_{E}$ is a quantum evaluation oracle for $X$ if
$\mathcal{U}_{X}:|\omega\rangle|0\rangle\mapsto|\omega\rangle|X(\omega)\rangle$
for all $\omega\in\Omega$, assuming $0\in E$.
@@
Next, we present a key quantum mean estimation result that will be carefully employed in our algorithmic framework to utilize the quantum superpositioned states collected by the RL agent. One of the crucial aspects of Lemma 1 is that quantum mean estimation converges at the rate $\mathcal{O}(\frac{1}{n})$ as opposed to classical benchmark convergence rate $\mathcal{O}(\frac{1}{\sqrt{n}})$ for $n$ number of samples, therefore estimation efficiency quadratically.
Lemma 1 (Quantum multivariate bounded estimator, Theorem 3.3 of (?)).
Let $X$ be a d-dimensional bounded random variable such that $||X||_{2}\leq 1$. Given three reals $L_{2}\in(0,1]$, $\delta\in(0,1)$ and $n\geq 1$ such that $\mathbb{E}[||X||_{2}]\leq L_{2}$, the multivariate bounded estimator $\texttt{QBounded}_{d}(X,L_{2},n,\delta)$ obtained by Algorithm 1 outputs a mean estimate $\hat{\mu}$ of $\mu=\mathbb{E}[X]$ such that,
${\mathbb{P}\left[||\hat{\mu}-\mu||_{\infty}\leq\frac{\sqrt{L_{2}}\log(d/\delta)}{n}\right]\geq 1-\delta.}$ (1)
The algorithm performs $\mathcal{O}\Big{(}n\log^{1/2}(n\sqrt{d})\Big{)}$ quantum experiments.
Algorithm 1 $\texttt{QBounded}_{d}(X,L_{2},n,\delta)$ (Algorithm 1 in (?))
1:if $n\leq\frac{\log(d/\delta)}{\sqrt{L_{2}}}$ then
2:     Output $\hat{\mu}=0$.
3:end if
@@
4:Set $\alpha=\frac{1}{\sqrt{\log(400\pi n\sqrt{d})}}$ and $m=2^{\left\lceil\log\left(\frac{8\pi n}{\alpha\sqrt{L_{2}\log(d/\delta)}}\right)\right\rceil}$.
5:Set $G=\left\{\frac{j}{m}-\frac{1}{2}+\frac{1}{2m}:j\in\{0,\dots,m-1\}\right\}^{d}\subseteq\left(-\frac{1}{2},\frac{1}{2}\right)^{d}$
6:for $k=1,\dots,\lceil 18\log(d/\delta)\rceil$ do
7:     Compute the uniform superposition $|G\rangle:=\frac{1}{\sqrt{md/2}}\sum_{u\in G}|u\rangle$ over $G$.
8:     Compute the state $|\psi\rangle:=\widetilde{P}_{X,L_{2},m,\alpha,\epsilon}|G\rangle$ in $H_{G}\otimes H_{aux}$, where $\widetilde{P}_{X,L_{2},m,\alpha,\epsilon}$ is the directional mean oracle constructed by the quantum evaluation oracle in Proposition 3.2 (?) with $\epsilon=1/25$.
9:     Compute the state $|\phi\rangle:=(QFT_{G}^{-1}\otimes I_{aux})|\psi\rangle$ where the unitary $QFT_{G}:|u\rangle\mapsto\frac{1}{\sqrt{md/2}}\sum_{v\in G}e^{2\pi i\langle u,v\rangle}|v\rangle$ is the quantum Fourier transform over $G$.
10:     Measure the $H_{G}$ register of $|\phi\rangle$ in the computational basis and let $v^{(k)}\in G$ denote the obtained result. Set $\mu^{(k)}=\frac{2\pi v^{(k)}}{\alpha}$.
11:end for
12:Output the coordinate-wise median $\hat{\mu}=\text{median}(\mu^{(1)},\dots,\mu^{(\lceil 18\log(d/\delta)\rceil)})$.
@@
Quantum Reinforcement Learning: Within the realm of QRL, a prominent strand of previous research showcases exponential enhancements in regret through amplitude amplification-based methodologies applied to the Quantum Multi-Armed Bandits (Q-MAB) problem (?, ?). Nevertheless, the theoretical underpinnings of the aforementioned Q-MAB approaches do not seamlessly extend to the QRL context since there is no state evolution in bandits.
@@
A recent surge of research interest has been directed towards Quantum Reinforcement Learning (QRL) (?, ?, ?). It is noteworthy to emphasize that the aforementioned studies do not provide an exhaustive mathematical characterization of regret. (?) delves into QRL within the infinite horizon Markov Decision Process (MDP) learning framework, yet their focus rests on a discounted reward setup, where advancements in sample complexity are demonstrated through the lens of generative model assumptions. Further, finite horizon episodic MDP setup has been considered in (?). Ours is the first paper to study infinite horizon MDP with average reward objective.
3 Problem Formulation
@@
We consider the problem of Reinforcement Learning (RL) in an infinite horizon Markov Decision Process characterized by $\mathcal{M}\triangleq(\mathcal{S},\mathcal{A},P,r,D)$, wherein $\mathcal{S}$ and $\mathcal{A}$ represent finite collections of states and actions respectively with $|\mathcal{S}|=S$ and $|\mathcal{A}|=A$, pertaining to RL agent’s interaction with the unknown MDP environment. $P(s^{\prime}|s,a)\in[0,1]$ denotes the transition probability for next state $s^{\prime}\in\mathcal{S}$ for a given pair of previous state and RL agent’s action, i.e., $(s,a)\in\mathcal{S}\times\mathcal{A}$. Further, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ represents the reward collected by the RL agent for state-action pair $(s,a)$. $D$ is the diameter of the MDP $\mathcal{M}$. In the following, we first present the additional quantum transition oracle that the agent could access during its interaction with the unknown MDP environment at every RL round. Here, we would like to emphasize that unknown MDP environment implies that matrix $P$ which encapsulates the transition dynamics of the underlying RL environment is not known beforehand.
@@
Quantum Computing Basics: In this section, we give a brief introduction to the concepts that are most related to our work based on (?) and (?). Consider an $m$ dimensional Hilbert space $\mathbb{C}^{m}$, a quantum state $|x\rangle=(x_{1},...,x_{m})^{T}$ can be seen as a vector inside the Hilbert space with $\sum_{i}|x_{i}|^{2}=1$. Furthermore, for a finite set with $m$ elements $\xi=\{\xi_{1},...,\xi_{m}\}$, we can assign each $\xi_{i}\in\xi$ with a member of an orthonormal basis of the Hilbert space $\mathbb{C}^{m}$ by
@@
Quantum transition oracle: We utilize the concepts and notations of quantum computing in (?, ?, ?, ?) to construct the quantum sampling oracle of RL environments and capture agent’s interaction with the unknown MDP environment. We now formulate the equivalent quantum-accessible RL environments for our classical MDP $\mathcal{M}$. For an agent at step $t$ in state $s_{t}$ and with action $a_{t}$, we construct the quantum sampling oracles for variables of the next state $P(\cdot|s_{t},a_{t})$. Specifically, suppose there are two Hilbert spaces $\mathcal{\bar{S}}=\mathbb{C}^{|\mathcal{S}|}$ and $\mathcal{\bar{A}}=\mathbb{C}^{|\mathcal{A}|}$ containing the superpositions of the classical states and actions. We represent the computational bases for $\mathcal{\bar{S}}$ and $\mathcal{\bar{A}}$ as $\{|s\rangle\}_{s\in\mathcal{S}}$ and $\{|a\rangle\}_{a\in\mathcal{A}}$. We assume that we can implement the quantum sampling oracle $\mathcal{U}_{P}$ of the MDP’s transitions as follows:
  • The quantum evaluation oracle for the transition probability (quantum transition oracle) $\mathcal{U}_{P}$ which at step $t$, returns the superposition over $s^{\prime}\in\mathcal{S}$ according to $P(s^{\prime}|s_{t},a_{t})$, the probability distribution of the next state given the current state $|s_{t}\rangle$ and action $|a_{t}\rangle$ is defined as:
    $\mathcal{U}_{P}:|s_{t}\rangle\otimes|a_{t}\rangle\otimes|0\rangle\rightarrow|\bm{\psi}\rangle^{t}$
    Where $|\bm{\psi}\rangle^{t}=|s_{t}\rangle\otimes|a_{t}\rangle\otimes\sum_{s^{\prime}\in\mathcal{S}}\sqrt{P(s^{\prime}|s_{t},a_{t})}|s^{\prime}\rangle$.
@@
As described in Fig 1, at step $t$, the agent plays an action $a_{t}$ based on current state $s_{t}$ according to its policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$. Consequently, the unknown MDP Q-environment shares the next state information and reward i.e., $\{s_{t+1},r_{t}\}$. Additionally, the quantum transition oracle first encodes $s_{t}$ and $a_{t}$ into the corresponding basis of the quantum Hilbert spaces, producing basis $|s_{t}\rangle$ and $|a_{t}\rangle$, then it encodes the superpositioned quantum state $|\bm{\psi}\rangle^{t}$ from the quantum transition oracle $\mathcal{U}_{P}$. $|\bm{\psi}\rangle^{t}$ would be used for the Quantum Mean Estimator QBounded outlined in Lemma 1 which in turn improves transition probability estimation and leads to exponentially better regret. We note that a single copy of $|\bm{\psi}\rangle^{t}$ is obtained at each $(s_{t},a_{t})$. Given that a measurement collapses the state, we can only perform one measurement on it. Thus, the quantum mean estimator will only perform measurements at the end of epochs, as will be described in the algorithm.
RL regret formulation: Consider the RL agent following some policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ on the unknown MDP Q-environment $\mathcal{M}$. Then, as a result of playing policy $\pi$, the agent’s long term average reward is defined next according to (?).
@@
Then, according to (?), $\Gamma_{\pi}^{P}$ can be represented as the following:
@@
wherein, $V_{\pi}^{P}(\cdot;\cdot)$ is the discounted cumulative average reward by following policy $\pi$, $\rho_{\pi}^{P}$ is the steady-state occupancy measure (i.e., $\rho_{\pi}^{P}(s,a)\in[0,1],~{}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\rho_{\pi}^{P}(s,a)=1$), $\bar{r}(s,a)$ is the steady-state long-term average reward of pair $(s,a)$. In the following, we introduce key assumptions crucial from the perspective of our algorithmic framework and its analysis. In this context, we introduce the notations $\{P_{\pi,s}^{t},T_{\pi,~{}s\rightarrow s^{\prime}}\}$ to denote the $t$ time step probability distribution obtained by playing policy $\pi$ on the unknown MDP $\mathcal{M}$ starting from some arbitrary state $s\in\mathcal{S}$; and, the actual time steps to reach $s^{\prime}$ from $s$ by playing policy $\pi$ respectively. With these notations, we introduce our first assumption pertaining to ergodicity of MDP $\mathcal{M}$.
@@
We emphasize that Assumption 2 is a commonly used assumption in RL algorithm development owing to the fact that in most setups rewards are constructed according to the underlying problem and is known beforehand (?, ?). Next, we define the cumulative regret accumulated by the agent in expectation across $T$ time steps as follows: